One particular brand of bread has two different sizes of loaves; 9 inches in length and 12 inches in length. If the 9 in. loaf has 15 slices, how many slices should the longer loaf have? (Assume that the slices in each are the same thickness.)
Please show detailed steps on this. Thank you.

Answer by stanbon(75887)

One particular brand of bread has two different sizes of loaves; 9 inches in length and 12 inches in length. If the 9 in. loaf has 15 slices, how many slices should the longer loaf have?
---
Use a proportion:
slices/lenght = slices/length
x/12 in= 15/8
----
x = 12(15/9)
-----
x = 20 slices for 12 in loaf
================
Cheers,
Stan H.
======================